OpenStudy (anonymous):

Find and factor the derivative of the function f(x)= log8 (4^x +3^-x)

OpenStudy (anonymous):

I am always confused when I read log8, is log8 = lg8 or ln8? lg has a base of 10, ln has a base of e

OpenStudy (anonymous):

I am asking this because: \[\Large \log_8(4^x-3^{-x}) \] Would not be the same problem as \[ \Large \underbrace{\ln8}_{\text{a number}}(4^x-3^{-x}) \]

OpenStudy (anonymous):

he probably means log base 8. log being implied as ln is more of an advanced technique of physicists and chemists. Mathematicians wouldn't do such a thing.

OpenStudy (anonymous):

I agree @GregTheo. I keep asking nowadays because I have solved - and on top of that, caused a lot of confusion - by copying the wrong problem (-:

OpenStudy (anonymous):

@pdube Is it the logarithm base eight? \[\Large \log_8x \]

OpenStudy (anonymous):

yes ...sorry I took awhile to reply I lost connection

OpenStudy (anonymous):

I think I got it but I'm not sure if I got the derivative right

OpenStudy (anonymous):

No problem, well clearly what you have to do first is to use change of bases \[\Large \log_8x= \frac{\ln x}{\ln8} \] So you can substitute that back in for your function and you will get: \[\Large \frac{1}{\ln8} \cdot \ln (4^x-3^{-x})\]